OpenStudy (anonymous):

Solve the equation for x and round your answer to the nearest hundreth. 5 = ln(x + 1)

OpenStudy (mathstudent55):

Use the definition of ln. \( \ln x = y\) means that \(e^y = x\).

OpenStudy (anonymous):

I don't know what it is!

OpenStudy (anonymous):

I've tried

OpenStudy (mathstudent55):

Use the definition of ln, if \(\ln x = y\), then \(e^y = x\) with your problem. \( 5 = \ln(x + 1) \) \( \ln (x + 1) = 5\) means \(e^5 = x + 1\). \(e^5\) is just a number, so solve the equation \(e^5 = x + 1\) for x. (Hint: subtract 1 from both sides.)